# Tidal effects based on GUP-induced effective metric

### v1

@@ Abstract
Quantum gravity phenomenology predicts the possible existence of a minimal length on the smallest scale Maggiore:1993rv ; Kempf:1994su , which suggests that the Heisenberg uncertainty principle (HUP) in quantum mechanics should be modified to use a generalized uncertainty principle (GUP) Garay:1994en ; Scardigli:1999jh ; KalyanaRama:2001xd ; Chang:2001bm ; Hossenfelder:2012jw ; Tawfik:2014zca . In this line of research, various types of GUPs Bolen:2004sq ; Bambi:2007ty ; Park:2007az ; Pedram:2011gw ; Nozari:2012gd ; Chung:2019raj have been studied with much interest over the last few decades to extend the current understanding of general relativity (GR) to the quantum gravity regime. In particular, using the HUP and GUPs, quantum effects on classical GR have been explored through the study of Hawking radiation Hawking:1974sw , the thermal radiation emitted by a black hole. The results heuristically include the Hawking temperature. Having obtained the Hawking temperature, one can proceed to explore black hole thermodynamics Adler:2001vs ; Nozari:2005ah ; Xiang:2006ei ; Myung:2006qr ; Hai-Xia:2007 ; Xiang:2009yq ; Banerjee:2010sd ; Gangopadhyay:2013ofa ; Gine:2020ves ; Lutfuoglu:2021ofc ; Su:2022cwd ; Yu:2024qzl and further to statistical mechanics Liu:2001ra ; Li:2002xb ; Kim:2007nh ; Nouicer:2007jg ; Kim:2007if ; Zhao:2009zzb ; Tang:2017wph ; Hong:2021xeg .
However, as is well known, Einstein’s GR is the geometric theory of gravitation and a given spacetime structure is completely determined by a metric tensor as a solution to Einstein’s equations, which describe the relation between the geometry of a spacetime and the energy-momentum contained in that spacetime. In this regard, if one can find a metric tensor including the GUP effect, it would be more efficient to implement research such as black hole physics, including both classical and quantum aspects. Various authors Scardigli:2014qka ; FaragAli:2015boi ; Contreras:2016xib ; Vagenas:2017vsw ; Jusufi:2022uhk ; Chemisana:2023fuk have tried to incorporate GUPs into metrics, which we will call a GUP-induced effective metric. Very recently, Ong Ong:2023jkp has obtained one by careful reasoning and calculated physical quantities like Hawking temperature and black hole shadow.
@@ II GUP-induced effective metric
Hereafter, we will use the units of $\hbar=L_{p}=1$ unless otherwise specified.
@@
$T_{\mathrm{GUP}}=\frac{M}{\pi\alpha}\left(1-\sqrt{1-\frac{\alpha}{4M^{2}}}\right).$ (2.4)
Here, we have followed the convention of introducing $1/2\pi$ factor to give the Hawking temperature of the Schwarzschild black hole, $T_{\rm Sch}=1/8\pi M$, in the small $\alpha$ limit.
@@
$f(r)=\left(1-\frac{2M}{r}\right)g(r),$ (2.6)
@@
$T=\frac{1}{4\pi}\left.f^{\prime}(r)\right|_{r=r_{H}}=\frac{g(r_{H})}{8\pi M},$ (2.7)
proportional to $g(r_{H})$ function. Note that the event horizon remains intact, i.e., at $r_{H}=2M$, $f(r_{H})=0$ although $g(r_{H})\neq 0$.
@@
$g(r_{H})=\frac{2r^{2}_{H}}{\alpha}\left(1-\sqrt{1-\frac{\alpha}{r^{2}_{H}}}\right).$ (2.8)
Thus, one can infer the proper form of $g(r)$ as
$g(r)=\frac{2r^{2}}{\alpha}\left(1-\sqrt{1-\frac{\alpha}{r^{2}}}\right),$ (2.9)
@@ III Geodesic in Schwarzschild black hole with effective metric
$\frac{d^{2}x^{\mu}}{d\tau^{2}}+\Gamma^{\mu}_{\nu\rho}\frac{dx^{\nu}}{d\tau}\frac{dx^{\rho}}{d\tau}=0,$ (3.15)
@@
$\frac{dv^{0}}{d\tau}-\frac{\left(1-\frac{2M}{r}-\sqrt{1-\frac{\alpha}{r^{2}}}\right)}{r\left(1-\frac{2M}{r}\right)\sqrt{1-\frac{\alpha}{r^{2}}}}v^{0}v^{1}=0,$ (3.17)
$\frac{dv^{1}}{d\tau}+\frac{2r\left(1-\frac{2M}{r}\right)\left(1-\sqrt{1-\frac{\alpha}{r^{2}}}\right)\left(\alpha-2r^{2}\left(1-\frac{M}{r}\right)\left(1-\sqrt{1-\frac{\alpha}{r^{2}}}\right)\right)}{\alpha^{2}\sqrt{1-\frac{\alpha}{r^{2}}}}(v^{0})^{2}+\frac{\left(1-\frac{2M}{r}-\sqrt{1-\frac{\alpha}{r^{2}}}\right)}{2r\left(1-\frac{2M}{r}\right)\sqrt{1-\frac{\alpha}{r^{2}}}}(v^{1})^{2}$
@@
$\frac{dv^{2}}{d\tau}+\frac{2}{r}v^{1}v^{2}-\sin\theta\cos\theta(v^{3})^{2}=0,$ (3.19)
$\frac{dv^{3}}{d\tau}+\frac{2}{r}v^{1}v^{3}+2\cot\theta v^{2}v^{3}=0,$ (3.20)
where we denote the four-velocity vector as $v^{\mu}=dx^{\mu}/d\tau$. For simplicity, one can consider the geodesics on the equatorial plane $\theta=\pi/2$ and thus $v^{2}=\dot{\theta}=0$ for all $\tau$, without loss of generality. Then, the geodesic equations are simplified to
$\frac{dv^{0}}{d\tau}-\frac{\left(1-\frac{2M}{r}-\sqrt{1-\frac{\alpha}{r^{2}}}\right)}{r\left(1-\frac{2M}{r}\right)\sqrt{1-\frac{\alpha}{r^{2}}}}v^{0}v^{1}=0,$ (3.21)
$\frac{dv^{1}}{d\tau}+\frac{2r\left(1-\frac{2M}{r}\right)\left(1-\sqrt{1-\frac{\alpha}{r^{2}}}\right)\left(\alpha-2r^{2}\left(1-\frac{M}{r}\right)\left(1-\sqrt{1-\frac{\alpha}{r^{2}}}\right)\right)}{\alpha^{2}\sqrt{1-\frac{\alpha}{r^{2}}}}(v^{0})^{2}+\frac{\left(1-\frac{2M}{r}-\sqrt{1-\frac{\alpha}{r^{2}}}\right)}{2r\left(1-\frac{2M}{r}\right)\sqrt{1-\frac{\alpha}{r^{2}}}}(v^{1})^{2}$
@@
can be used to identify the integration constants $c_{1}$, $c_{2}$ with $E$, $L$, respectively. Here, $\xi^{\mu}=(1,0,0,0)$ and $\psi^{\mu}=(0,0,0,1)$ are the Killing vectors. Finally, by letting $ds^{2}=-kd\tau^{2}$ and using Eqs. (3.10) and (3.11), one can obtain
$v^{1}=\frac{dr}{d\tau}=\pm\left[E^{2}-\left(k+\frac{L^{2}}{r^{2}}\right)\left(1-\frac{2M}{r}\right)\frac{2r^{2}}{\alpha}\left(1-\sqrt{1-\frac{\alpha}{r^{2}}}\right)\right]^{1/2},$ (3.28)
@@ IV Tidal force in Schwarzschild black hole with effective metric
$\frac{D^{2}\eta^{\mu}}{D\tau^{2}}+R^{\mu}_{\nu\rho\sigma}v^{\nu}\eta^{\rho}v^{\sigma}=0,$ (4.1)
@@
$\frac{d^{2}\eta^{\hat{1}}}{d\tau^{2}}$ $=$ $-\frac{f^{\prime\prime}(r)}{2}\eta^{\hat{1}}=\frac{(2M-3r)\alpha+2r^{3}\left(1-\left(1-\frac{\alpha}{r^{2}}\right)^{3/2}\right)}{r^{3}\alpha\left(1-\frac{\alpha}{r^{2}}\right)^{3/2}}\eta^{\hat{1}},$ (4.5)
$\frac{d^{2}\eta^{\hat{i}}}{d\tau^{2}}$ $=$ $-\frac{f^{\prime}(r)}{2r}\eta^{\hat{i}}=-\frac{\alpha+2(M-r)r\left(1-\sqrt{1-\frac{\alpha}{r^{2}}}\right)}{r^{2}\alpha\sqrt{1-\frac{\alpha}{r^{2}}}}\eta^{\hat{i}},$ (4.6)
@@
$\frac{d^{2}\eta^{\hat{1}}}{d\tau^{2}}$ $=$ $\frac{2M}{r^{3}}\eta^{\hat{1}},$ (4.7)
$\frac{d^{2}\eta^{\hat{i}}}{d\tau^{2}}$ $=$ $-\frac{M}{r^{3}}\eta^{\hat{i}},$ (4.8)
@@ V Geodesic deviation equations of Schwarzschild black hole with effective metric
$[E^{2}-f(r)]\frac{d^{2}\eta^{\hat{1}}}{dr^{2}}-\frac{f^{\prime}(r)}{2}\frac{d\eta^{\hat{1}}}{dr}+\frac{f^{\prime\prime}(r)}{2}\eta^{\hat{1}}$ $=$ $0,$ (5.1)
$[E^{2}-f(r)]\frac{d^{2}\eta^{\hat{i}}}{dr^{2}}-\frac{f^{\prime}(r)}{2}\frac{d\eta^{\hat{i}}}{dr}+\frac{f^{\prime}(r)}{2r}\eta^{\hat{i}}$ $=$ $0$ (5.2)
@@
$\eta^{\hat{1}}(r)=c_{1}\sqrt{E^{2}-f(r)}+c_{2}\sqrt{E^{2}-f(r)}\int\frac{dr}{[E^{2}-f(r)]^{3/2}},$ (5.3)
@@
$\eta^{\hat{i}}(r)=r\left(c_{3}+c_{4}\int\frac{dr}{r^{2}\sqrt{E^{2}-f(r)}}\right),$ (5.4)
@@
Figure 5: Radial components of geodesic deviation in the Schwarzschild black hole modified by an effective metric, (a) with $\frac{d\eta^{\hat{1}}(b)}{d\tau}=0$ and (b) with $\frac{d\eta^{\hat{1}}(b)}{d\tau}=1~{}(\neq 0)$, the dashed curves are for $\alpha=0$, and the solid curves for $\alpha=1.0,~{}2.0,~{}3.0$ from the left.
Figure 6: Radial components of geodesic deviation by varying the mass $M$ in the Schwarzschild black hole modified by an effective metric with $\frac{d\eta^{\hat{1}}(b)}{d\tau}=1$. The dashed curve is for $M=1$, and the solid curves are for $M=5,~{}10,~{}15$ downwards.
@@
Note that $d_{2}(r)$ has no $r$-dependence. The integration constants $c_{i}$ can be determined by adopting appropriate boundary conditions that the infinitesimal displacement and initial velocity between two nearby particles are given by $\eta^{k}(b)$ and $d\eta^{k}(r)/d\tau|_{r=b}\equiv d\eta^{k}(b)/d\tau$ ($k=1,~{}2,~{}3$) at $r=b$. Explicitly, we have
$c_{1}$ $=$ $\frac{b^{2}}{M}\left(1+\alpha Q^{(1)}(b)+\alpha^{2}Q^{(2)}(b)\right)^{-1}\frac{d\eta^{\hat{1}}(b)}{d\tau},$
@@
$c_{4}$ $=$ $-b\left(1+\alpha Q^{(1)}(b)+\alpha^{2}Q^{(2)}(b)\right)^{-1/2}d^{-1}_{3}(b)\frac{d\eta^{\hat{i}}(b)}{d\tau}.$ (5.10)
@@
$\eta^{\hat{1}}(r)$ $=$ $b\sqrt{\frac{2b}{M}}\frac{d\eta^{\hat{1}}(b)}{d\tau}\frac{\sqrt{br-r^{2}}}{r}\frac{\left(1+\alpha Q^{(1)}(r)+\alpha^{2}Q^{(2)}(r)\right)^{1/2}}{\left(1+\alpha Q^{(1)}(b)+\alpha^{2}Q^{(2)}(b)\right)}$ (5.11)
@@
$\eta^{\hat{i}}(r)$ $=$ $\frac{1}{b}\eta^{\hat{i}}(b)r+\sqrt{\frac{2b}{M}}\frac{d\eta^{\hat{i}}(b)}{d\tau}d^{-1}_{2}(b)\left(\frac{br-r^{2}}{1+\alpha Q^{(1)}(b)+\alpha^{2}Q^{(2)}(b)}\right)^{1/2}d_{3}(r),$ (5.12)
Figure 7: Angular components of geodesic deviation in the Schwarzschild black hole modified by the $\alpha$ dependent effective metric, (a) with $\alpha=0$ and $\frac{d\eta^{\hat{1}}(b)}{d\tau}=0$, (b) with $\alpha=0$ and $\frac{d\eta^{\hat{1}}(b)}{d\tau}=1$ for the dashed curve, and $\alpha=1.0,~{}2.0,~{}3.0$ (from the left near $r=0.5$) and $\frac{d\eta^{\hat{1}}(b)}{d\tau}=1$ for the solid curves.
@@
In Fig. 5, we have drawn the radial components of the geodesic deviation by comparing the case of not having GUP parameter to that of having a GUP parameter for both $\frac{d\eta^{\hat{1}}(b)}{d\tau}=0$ and $\frac{d\eta^{\hat{1}}(b)}{d\tau}\neq 0$ initial velocities. From the figure, one can see that the radial separation vectors of a falling object in the GUP-modified Schwarzschild black hole start with the same initial values as the original Schwarzschild black hole, and then it gets stretched steeper than the original Schwarzschild case as it approaches the singularities. This results in much greater separation effects. Note that comparing Fig. 5(a) with Fig. 5(b), the latter shows a much steeper slope in the range of $100<r<$ 10 than the former. In Fig. 6, we have also drawn the radial components of the geodesic deviation by varying the mass parameter $M$ in the Schwarzschild black hole modified by the effective metric. Here, one can find that the bigger the black hole mass is, the greater the geodesic separation is near the singularity.

### v2

@@ Abstract
Quantum gravity phenomenology predicts the possible existence of a minimal length on the smallest scale Maggiore:1993rv ; Kempf:1994su , which suggests that the Heisenberg uncertainty principle (HUP) in quantum mechanics should be modified to use a generalized uncertainty principle (GUP) Garay:1994en ; Scardigli:1999jh ; KalyanaRama:2001xd ; Chang:2001bm ; Hossenfelder:2012jw ; Tawfik:2014zca . In this line of research, various types of GUPs Bolen:2004sq ; Bambi:2007ty ; Park:2007az ; Pedram:2011gw ; Nozari:2012gd ; Chung:2019raj have been studied with much interest over the last few decades to extend the current understanding of general relativity (GR) to the quantum gravity regime. In particular, using the HUP and GUPs, quantum effects on classical GR have been explored through the study of Hawking radiation Hawking:1974sw , the thermal radiation emitted by a black hole. The results heuristically include the Hawking temperature. Having obtained the Hawking temperature, one can proceed to study black holes Adler:2001vs ; Nozari:2005ah ; Medved:2004yu ; Chen:2014jwq ; Carr:2015nqa ; Bosso:2023aht including their thermodynamics Amelino-Camelia:2005zpp ; Xiang:2006ei ; Myung:2006qr ; Hai-Xia:2007 ; Xiang:2009yq ; Banerjee:2010sd ; Gangopadhyay:2013ofa ; Gine:2020ves ; Lutfuoglu:2021ofc ; Su:2022cwd ; Yu:2024qzl and further to statistical mechanics Liu:2001ra ; Li:2002xb ; Kim:2007nh ; Nouicer:2007jg ; Kim:2007if ; Zhao:2009zzb ; Tang:2017wph ; Hong:2021xeg .
However, as is well known, Einstein’s GR is the geometric theory of gravitation and a given spacetime structure is completely determined by a metric tensor as a solution to Einstein’s equations, which describe the relation between the geometry of a spacetime and the energy-momentum contained in that spacetime. In this regard, if one can find a metric tensor including the GUP effect, it would be more efficient to implement research such as black hole physics, including both classical and quantum aspects. Various authors Scardigli:2014qka ; FaragAli:2015boi ; Vagenas:2017vsw ; Contreras:2016xib ; Anacleto:2020lel ; Anacleto:2021qoe ; Jusufi:2022uhk ; Chemisana:2023fuk have tried to incorporate GUPs into metrics, which we will call a GUP-induced effective metric. Very recently, Ong Ong:2023jkp has obtained a satisfying GUP-induced effective metric by emphasizing the importance of using the full GUP expression instead of just series expansions Contreras:2016xib ; Anacleto:2020lel ; Anacleto:2021qoe , which has some subtleties and shortcomings of undesired features.
@@ II GUP-induced effective metric
$T_{\mathrm{GUP}}=\frac{Mc}{\pi\alpha}\left(1-\sqrt{1-\frac{\alpha\hbar c}{4GM^{2}}}\right).$ (2.4)
Here, we have followed the convention of introducing $1/2\pi$ factor to give the Hawking temperature of the Schwarzschild black hole, $T_{\rm Sch}=1/8\pi M$, in the small $\alpha$ limit. Note also that $M^{2}_{p}=\hbar c/G$.
@@
$f(r)=\left(1-\frac{2GM}{rc^{2}}\right)g(r),$ (2.6)
@@
$T=\frac{1}{4\pi}\left.f^{\prime}(r)\right|_{r=r_{H}}=\frac{c^{2}}{8\pi GM}g(r_{H}),$ (2.7)
proportional to $g(r_{H})$ function. Note that the event horizon remains intact, i.e., at $r_{H}=2GM/c^{2}$, $f(r_{H})=0$ although $g(r_{H})\neq 0$.
@@
$g(r_{H})=\frac{2r^{2}_{H}c}{G\alpha}\left(1-\sqrt{1-\frac{\alpha L^{2}_{p}}{r^{2}_{H}}}\right),$ (2.8)
where $L^{2}_{p}=G\hbar/c^{3}$. Thus, one can infer the proper form of $g(r)$ as
$g(r)=\frac{2r^{2}c}{G\alpha}\left(1-\sqrt{1-\frac{\alpha L^{2}_{p}}{r^{2}}}\right),$ (2.9)
@@
$f(r)=\left(1-\frac{2GM}{rc^{2}}\right)\frac{2r^{2}c}{G\alpha}\left(1-\sqrt{1-\frac{\alpha L^{2}_{p}}{r^{2}}}\right).$ (2.10)
For the issue of uniqueness of choosing $g(r)$, we refer Ong’s work Ong:2023jkp . Hereafter, for simplicity, we will use the Planck units of $c=\hbar=G=1$, which also imply that $L_{p}=M_{p}=1$. Thus, the GUP-induced effective metric which we will study has the form of
@@ III Geodesic in Schwarzschild black hole with effective metric
$\frac{\rm{d}^{2}x^{\mu}}{\rm{d}\tau^{2}}+\Gamma^{\mu}_{\nu\rho}\frac{\rm{d}x^{\nu}}{\rm{d}\tau}\frac{\rm{d}x^{\rho}}{\rm{d}\tau}=0,$ (3.16)
@@
$\frac{\rm{d}v^{0}}{\rm{d}\tau}-\frac{\left(1-\frac{2M}{r}-\sqrt{1-\frac{\alpha}{r^{2}}}\right)}{r\left(1-\frac{2M}{r}\right)\sqrt{1-\frac{\alpha}{r^{2}}}}v^{0}v^{1}=0,$ (3.18)
$\frac{\rm{d}v^{1}}{\rm{d}\tau}+\frac{2r\left(1-\frac{2M}{r}\right)\left(1-\sqrt{1-\frac{\alpha}{r^{2}}}\right)\left(\alpha-2r^{2}\left(1-\frac{M}{r}\right)\left(1-\sqrt{1-\frac{\alpha}{r^{2}}}\right)\right)}{\alpha^{2}\sqrt{1-\frac{\alpha}{r^{2}}}}(v^{0})^{2}+\frac{\left(1-\frac{2M}{r}-\sqrt{1-\frac{\alpha}{r^{2}}}\right)}{2r\left(1-\frac{2M}{r}\right)\sqrt{1-\frac{\alpha}{r^{2}}}}(v^{1})^{2}$
@@
$\frac{\rm{d}v^{2}}{\rm{d}\tau}+\frac{2}{r}v^{1}v^{2}-\sin\theta\cos\theta(v^{3})^{2}=0,$ (3.20)
$\frac{\rm{d}v^{3}}{\rm{d}\tau}+\frac{2}{r}v^{1}v^{3}+2\cot\theta v^{2}v^{3}=0,$ (3.21)
where we denote the four-velocity vector as $v^{\mu}=\rm{d}x^{\mu}/\rm{d}\tau$. For simplicity, one can consider the geodesics on the equatorial plane $\theta=\pi/2$ and thus $v^{2}=\dot{\theta}=0$ for all $\tau$, without loss of generality. Then, the geodesic equations are simplified to
$\frac{\rm{d}v^{0}}{\rm{d}\tau}-\frac{\left(1-\frac{2M}{r}-\sqrt{1-\frac{\alpha}{r^{2}}}\right)}{r\left(1-\frac{2M}{r}\right)\sqrt{1-\frac{\alpha}{r^{2}}}}v^{0}v^{1}=0,$ (3.22)
$\frac{\rm{d}v^{1}}{\rm{d}\tau}+\frac{2r\left(1-\frac{2M}{r}\right)\left(1-\sqrt{1-\frac{\alpha}{r^{2}}}\right)\left(\alpha-2r^{2}\left(1-\frac{M}{r}\right)\left(1-\sqrt{1-\frac{\alpha}{r^{2}}}\right)\right)}{\alpha^{2}\sqrt{1-\frac{\alpha}{r^{2}}}}(v^{0})^{2}+\frac{\left(1-\frac{2M}{r}-\sqrt{1-\frac{\alpha}{r^{2}}}\right)}{2r\left(1-\frac{2M}{r}\right)\sqrt{1-\frac{\alpha}{r^{2}}}}(v^{1})^{2}$
@@
can be used to identify the integration constants $c_{1}$, $c_{2}$ with $E$, $L$, respectively. Here, $\xi^{\mu}=(1,0,0,0)$ and $\psi^{\mu}=(0,0,0,1)$ are the Killing vectors. Finally, by letting $\rm{d}s^{2}=-k\rm{d}\tau^{2}$ and using Eqs. (3.10) and (3.11), one can obtain
$v^{1}=\frac{\rm{d}r}{\rm{d}\tau}=\pm\left[E^{2}-\left(k+\frac{L^{2}}{r^{2}}\right)\left(1-\frac{2M}{r}\right)\frac{2r^{2}}{\alpha}\left(1-\sqrt{1-\frac{\alpha}{r^{2}}}\right)\right]^{1/2},$ (3.29)
@@ IV Tidal force in Schwarzschild black hole with effective metric
$\frac{\rm{D}^{2}\eta^{\mu}}{\rm{D}\tau^{2}}+R^{\mu}_{\nu\rho\sigma}v^{\nu}\eta^{\rho}v^{\sigma}=0,$ (4.1)
@@
$\frac{\rm{d}^{2}\eta^{\hat{1}}}{\rm{d}\tau^{2}}$ $=$ $-\frac{f^{\prime\prime}(r)}{2}\eta^{\hat{1}}=\frac{(2M-3r)\alpha+2r^{3}\left(1-\left(1-\frac{\alpha}{r^{2}}\right)^{3/2}\right)}{r^{3}\alpha\left(1-\frac{\alpha}{r^{2}}\right)^{3/2}}\eta^{\hat{1}},$ (4.5)
$\frac{\rm{d}^{2}\eta^{\hat{i}}}{\rm{d}\tau^{2}}$ $=$ $-\frac{f^{\prime}(r)}{2r}\eta^{\hat{i}}=-\frac{\alpha+2(M-r)r\left(1-\sqrt{1-\frac{\alpha}{r^{2}}}\right)}{r^{2}\alpha\sqrt{1-\frac{\alpha}{r^{2}}}}\eta^{\hat{i}},$ (4.6)
@@
$\frac{\rm{d}^{2}\eta^{\hat{1}}}{\rm{d}\tau^{2}}$ $=$ $\frac{2M}{r^{3}}\eta^{\hat{1}},$ (4.7)
$\frac{\rm{d}^{2}\eta^{\hat{i}}}{\rm{d}\tau^{2}}$ $=$ $-\frac{M}{r^{3}}\eta^{\hat{i}},$ (4.8)
@@ V Geodesic deviation equations of Schwarzschild black hole with effective metric
$[E^{2}-f(r)]\frac{\rm{d}^{2}\eta^{\hat{1}}}{\rm{d}r^{2}}-\frac{f^{\prime}(r)}{2}\frac{\rm{d}\eta^{\hat{1}}}{\rm{d}r}+\frac{f^{\prime\prime}(r)}{2}\eta^{\hat{1}}$ $=$ $0,$ (5.1)
$[E^{2}-f(r)]\frac{\rm{d}^{2}\eta^{\hat{i}}}{\rm{d}r^{2}}-\frac{f^{\prime}(r)}{2}\frac{\rm{d}\eta^{\hat{i}}}{\rm{d}r}+\frac{f^{\prime}(r)}{2r}\eta^{\hat{i}}$ $=$ $0$ (5.2)
@@
$\eta^{\hat{1}}(r)=c_{1}\sqrt{E^{2}-f(r)}+c_{2}\sqrt{E^{2}-f(r)}\int\frac{\rm{d}r}{[E^{2}-f(r)]^{3/2}},$ (5.3)
@@
$\eta^{\hat{i}}(r)=r\left(c_{3}+c_{4}\int\frac{\rm{d}r}{r^{2}\sqrt{E^{2}-f(r)}}\right),$ (5.4)
@@
Note that $d_{2}(r)$ has no $r$-dependence. The integration constants $c_{i}$ can be determined by adopting appropriate boundary conditions that the infinitesimal displacement and initial velocity between two nearby particles are given by $\eta^{k}(b)$ and $\rm{d}\eta^{k}(r)/\rm{d}\tau|_{r=b}\equiv\rm{d}\eta^{k}(b)/\rm{d}\tau$ ($k=1,~{}2,~{}3$) at $r=b$. Explicitly, we have
$c_{1}$ $=$ $\frac{b^{2}}{M}\left(1+\alpha Q^{(1)}(b)+\alpha^{2}Q^{(2)}(b)\right)^{-1}\frac{\rm{d}\eta^{\hat{1}}(b)}{\rm{d}\tau},$
@@
$c_{4}$ $=$ $-b\left(1+\alpha Q^{(1)}(b)+\alpha^{2}Q^{(2)}(b)\right)^{-1/2}d^{-1}_{3}(b)\frac{\rm{d}\eta^{\hat{i}}(b)}{\rm{d}\tau}.$ (5.10)
@@
$\eta^{\hat{1}}(r)$ $=$ $b\sqrt{\frac{2b}{M}}\frac{\rm{d}\eta^{\hat{1}}(b)}{\rm{d}\tau}\frac{\sqrt{br-r^{2}}}{r}\frac{\left(1+\alpha Q^{(1)}(r)+\alpha^{2}Q^{(2)}(r)\right)^{1/2}}{\left(1+\alpha Q^{(1)}(b)+\alpha^{2}Q^{(2)}(b)\right)}$ (5.11)
@@
$\eta^{\hat{i}}(r)$ $=$ $\frac{1}{b}\eta^{\hat{i}}(b)r+\sqrt{\frac{2b}{M}}\frac{\rm{d}\eta^{\hat{i}}(b)}{\rm{d}\tau}d^{-1}_{2}(b)\left(\frac{br-r^{2}}{1+\alpha Q^{(1)}(b)+\alpha^{2}Q^{(2)}(b)}\right)^{1/2}d_{3}(r),$ (5.12)
Figure 5: Radial components of geodesic deviation in the Schwarzschild black hole modified by an effective metric, (a) with $\frac{\rm{d}\eta^{\hat{1}}(b)}{\rm{d}\tau}=0$, and (b) with $\frac{\rm{d}\eta^{\hat{1}}(b)}{\rm{d}\tau}=1~{}(\neq 0)$, the dashed curves are for $\alpha=0$, and the solid curves for $\alpha=1.0,~{}2.0,~{}3.0$ from the left.
Figure 6: Radial components of geodesic deviation by varying the mass $M$ in the Schwarzschild black hole modified by an effective metric with $\frac{\rm{d}\eta^{\hat{1}}(b)}{\rm{d}\tau}=1$. (a) the solid curves for $M=5,~{}10,~{}15$, (b) the solid curves for $M=100,~{}200,~{}300$, downwards, with the dashed curve for $M=1$.
Figure 7: Angular components of geodesic deviation in the Schwarzschild black hole modified by the $\alpha$ dependent effective metric, (a) with $\alpha=0$ and $\frac{\rm{d}\eta^{\hat{1}}(b)}{\rm{d}\tau}=0$, (b) with $\alpha=0$ and $\frac{\rm{d}\eta^{\hat{1}}(b)}{\rm{d}\tau}=1$ for the dashed curve, and $\alpha=1.0,~{}2.0,~{}3.0$ (from the left near $r=0.5$) and $\frac{\rm{d}\eta^{\hat{1}}(b)}{\rm{d}\tau}=1$ for the solid curves.
@@
Figure 9: (a) Angular components of geodesic deviation by varying the mass $M$ in the effective metric for $M=1,~{}100,~{}200,~{}300$ (from top to down) with $\alpha=0$ in the original Schwarzschild black hole, (b) Angular components of geodesic deviation by varying the mass in the effective metric for $M=1,~{}100,~{}200,~{}300$ with $\alpha=1$ in the Schwarzschild black hole modified by an effective metric. Here, the dashed curve is for $\alpha=0$ case, as before, drawn for comparison.
@@
In Fig. 5, we have drawn the radial components of the geodesic deviation by comparing the case of not having GUP parameter to that of having a GUP parameter for both $\frac{\rm{d}\eta^{\hat{1}}(b)}{\rm{d}\tau}=0$ and $\frac{\rm{d}\eta^{\hat{1}}(b)}{\rm{d}\tau}\neq 0$ initial velocities. From the figure, one can see that the radial separation vectors of a falling object in the GUP-modified Schwarzschild black hole start with the same initial values as the original Schwarzschild black hole, and then it gets stretched steeper than the original Schwarzschild case as it approaches the singularities. This results in much greater separation effects. Note that comparing Fig. 5(a) with Fig. 5(b), the latter shows a much steeper slope in the range of $100<r<$ 10 than the former. In Fig. 6, we have also drawn the radial components of the geodesic deviation by varying the mass parameter $M$ in the Schwarzschild black hole modified by the effective metric. Here, one can find that the bigger the black hole mass is, the greater the geodesic separation is near the singularity. Moreover, for comparison, we have drawn the radial components of geodesic deviation for the Planck mass comparable scales $M$ in Fig. 6(a) and for some larger $M$ in Fig. 6(b), which shows that the latter sharply increases compared to the former as it approaches the singularity.